OpenStudy (megannicole51):

Find a formula for Sn, (n is greater than or equal to 1) 4,8,16,32,64 @agent0smith

OpenStudy (anonymous):

lol

OpenStudy (agent0smith):

Find the common ratio, divide one term by the term before it

OpenStudy (anonymous):

i'm not saying anything until i find the answer to last nights question

OpenStudy (agent0smith):

^lol she posted it @satellite73 in the question.

OpenStudy (anonymous):

btw this is a sequence, not a sum, maybe it is \(a_n\) rather than \(S_n\)

OpenStudy (megannicole51):

no its Sn.

OpenStudy (agent0smith):

Remember this from yesterday \[\Large S_n =a_1 \frac{ 1-r^n }{ 1-r }\]hopefully i remember that right.

OpenStudy (anonymous):

ooh i see! i mean the one from last night! how did you get it? did you read that business i sent, or did you figure it out some other way?

OpenStudy (megannicole51):

yes i remember that dang formula that i used all night lol

OpenStudy (megannicole51):

i figured it out another way.

OpenStudy (agent0smith):

:D well from 4,8,16,32,64 just find a1 and r.

OpenStudy (megannicole51):

A1 is 4 and r is 2?

OpenStudy (anonymous):

at least there was a 60.21 in the answer yes to your last question

OpenStudy (megannicole51):

yeah exactly...thanks again.

OpenStudy (anonymous):

ho ho you are welcome, even though it wasn't quite right you to this one now yes?

OpenStudy (anonymous):

* you got this one

OpenStudy (megannicole51):

so im just using the same formula as last night?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

if it was me, i would reverse the order top and bottom because \(r=2\) and \(2>1\) so i would write \[4\frac{2^n-1}{2-1}=4(2^n-1)\]

OpenStudy (agent0smith):

@satellite73 I like how you mentioned to a pretty smart calculus student that 2>1 :D

OpenStudy (megannicole51):

lol thanks....but why would you switch them? and are you able to do it?

OpenStudy (anonymous):

it is just if \(r>1\) you would have a negative number in the denominator (and also the numerator) since the denominator is \(1-r\)

OpenStudy (anonymous):

and "pretty smart calc student" or no, it is always the case that \[\frac{a-b}{c-d}=\frac{b-a}{d-c}\]

OpenStudy (agent0smith):

basically factoring out a negative \[\Large \frac{ a - b^n }{ a-b } = \frac{-( -a + b^n )}{ -( -a+b) }\]

OpenStudy (anonymous):

so if \(r>1\) you know two things: the series does not converge, i.e. you cannot add up an infinite number and also you probably want to use \[a_1(\frac{r^n-1}{r-1})\]

OpenStudy (anonymous):

especially if \(r=2\) because then \(r-1=1\)

OpenStudy (agent0smith):

@satellite73 i just don't think you found it funny like i did :P the 2>1 :D

OpenStudy (anonymous):

no i liked it

OpenStudy (megannicole51):

okay lets drop it...obviously im not a genius in math but im trying my best...

OpenStudy (anonymous):

here is a fact that you might want to commit to memory \[1+2+4+8+16+..=\sum_{k=0}^n2^k=2^{n+1}-1\] i think you are a math genius, i am just funnin please do not take offence

OpenStudy (megannicole51):

so thats my answer? and sorry i dont really pick up on people being sarcastic when its over like facebook or text or open study...

OpenStudy (agent0smith):

\[\Large 4(2^n-1)\] this is the most simplified answer i think